# A time-stepping deep gradient flow method for option pricing∑in (rough) diffusion models

### v1

@@ 1. Introduction
Deep learning and neural network methods have also been applied in the field of rough stochastic volatility modeling. Horvath et al. [35] presented a neural network-based method that performs the calibration task for the full implied volatility surface in (rough) volatility models. Jacquier and Oumgari [36] introduced a deep learning–based algorithm to evaluate options in affine rough stochastic volatility models. Jacquier and Zuric [38] constructed a deep learning-based numerical algorithm to solve path-dependent partial differential equations (PDEs) arising in the context of rough volatility.
Many other methods have been utilized and applied to rough stochastic volatility models. Let us mention here, among others, that El Euch and Rosenbaum [22] have computed the characteristic function of the log‐price in the rough Heston model, thus opening the door for the application of Fourier transform methods for option pricing, Bayer et al. [11] proposed the application of adaptive sparse grids and quasi Monte Carlo methods for option pricing in the rough Bergomi model, while Bayer et al. [12] have applied Hairer’s theory of regularity structures in order to analyze rough volatility models. In addition, Bayer et al. [10], Horvath et al. [34] and Jacquier and Pannier [37], among others, have studied asymptotic properties of rough volatility models in various regimes.
The connection between PDEs and option pricing is deep and well-studied in the literature. Several methods have been developed in order to price options in diffusion or stochastic volatility models, especially in high-dimensional settings, using sparse grids, asymptotic expansions, finite difference and finite elements methods; see e.g. Reisinger and Wittum [44], Düring and Fournié [20], Hilber et al. [32] and the comprehensive book by Hilber et al. [33].
@@
The starting point for our research is the existence of a Markovian approximation for a rough volatility model or, more generally, another Markovian model that describes empirical data well; see e.g. Rømer [45] and Abi Jaber and Li [4]. Inspired by Georgoulis et al. [28], we consider the option pricing PDE, introduce a time-stepping discretization of the time derivative, and reformulate the spatial derivatives as an energy minimization problem; the main difference with Georgoulis et al. [28] is the presence of the drift (first order) term, which we treat in an explicit fashion in the discretization. This gradient flow formulation of the PDE gives rise to a “natural” cost function for the training, via stochastic gradient descent, of deep neural networks that approximate the solution of the PDE. The main advantages of this method are twofold: on the one hand, the neural network from the previous time step is a good initial guess for the neural network optimization in the current time step, thus reducing the number of training stages necessary. On the other hand, the energy formulation allows to consider only first-order derivatives, thus reducing the training time in each time step / training stage significantly. In addition, we utilize available information about the qualitative behavior of option prices in order to assist the training of the neural network. More specifically, the architecture of the neural network respects the asymptotic behavior of option prices for large levels of moneyness, and adheres to a priori known bounds for option prices. The empirical results show that the proposed method is accurate, with a small overall error in the order of $10^{-3}$ and comparable with other deep learning methods, while the training times are significantly small compared to other methods and do not grow significantly with the dimension of the underlying space.
@@ 2. Problem formulation
$\begin{cases}u_{t}(t,x)+\mathcal{A}u(t,x)+ru(t,x)=0,\quad&(t,x)\in[0,T]\times\Omega,\\ u(0,x)=\Psi(x),\quad&x\in\Omega,\end{cases}$ (2.1)
@@ 3. Energy minimization and a deep gradient flow method
Let us now explain how we will develop our deep neural network method for solving the PDE (2.1). We proceed in three steps. First, we discretize the time derivative in Subsection 3.1 and introduce a time-stepping scheme. Second, we rewrite the solution of the PDE as an energy minimization problem in Subsection 3.2, which gives rise to a “natural” cost function. Third, we approximate the minimizer by a neural network, which is trained using stochastic gradient descent, in Subsection 3.3
@@ 3.1. Time discretization
Let us divide the time interval $[0,T]$ into $N$ equally spaced intervals $(t_{k-1},t_{k}]$, with $h=t_{k}-t_{k-1}=\frac{1}{N}$ for $k=0,1,\dots,N$. Let $U^{k}$ denote the approximation to the solution of the PDE $u(t_{k})$ at time step $t_{k}$, using the following time discretization scheme
@@
where $F(u)=\mathbf{b}\cdot\nabla u$. Convergence analysis of this scheme appears in Akrivis and Smyrlis [5].
@@ 3.2. Variational formulation
We would like now to associate the discretized PDE in (3.1) with an energy functional, in order to define later a “natural” cost function for the training of the neural network. In other words, we would like to find an energy functional $I(u)$ such that $U^{k}$ is a critical point of $I$. Let us rewrite (3.1) as follows:
@@
$U^{k}$ solves (3.2) if and only if $U^{k}$ minimizes
@@
Let $v$ be a smooth function. Consider the function
@@
$\quad+\left.h\int_{\Omega}\frac{1}{2}\left(\left(\nabla v\right)^{\mathsf{T}}A\nabla\left(U^{k}+\tau v\right)+\left(\nabla U^{k}+\tau v\right)^{\mathsf{T}}A\nabla v\right)+r\left(U^{k}+\tau v\right)v+F\left(U^{k-1}\right)v\mathrm{d}x\right]_{\tau=0}$
@@
$=\int_{\Omega}\left(U^{k}-U^{k-1}\right)v+h\left(A\nabla U^{k}\cdot\nabla v+rU^{k}+F\left(U^{k-1}\right)\right)v\mathrm{d}x$
@@
where in the second to last equality we used that $A$ is symmetric, and in the last equality we used the divergence theorem. This equality holds for all $v$ if and only if (3.2) holds. ∎
@@ 3.3. Neural network approximation
$L^{k}\left(\theta;\mathbf{x}\right)$ $=\frac{|\Omega|}{2M}\sum_{i=1}^{M}\left(f^{k}(\mathbf{x}_{i};\theta)-f^{k-1}(\mathbf{x}_{i})\right)^{2}+hN^{k}\left(\theta;\mathbf{x}\right),$
@@
In order to minimize this cost function, we use the stochastic gradient descent algorithm, i.e. the iterative scheme:
$\theta_{n+1}=\theta_{n}-\alpha_{n}\nabla_{\theta}L^{k}(\theta_{n};\mathbf{x}).$
The hyper-parameter $\alpha_{n}$ is the step-size of our update, called the learning rate. An overview of the Time Deep Gradient Flow (TDGF) method appears in Algorithm 1.
@@ 4. Examples
Let us now outline the different models to which we will apply our method, and derive the appropriate form of the generator in the option pricing PDE for each of these models. In particular, we will treat the Black–Scholes model in Subsection 4.1, the Heston model in Subsection 4.2, and the lifted Heston model in Subsection 4.3.
@@ 4.3. Lifted Heston model
$=f_{S}rS\mathrm{d}t-\sum_{i=1}^{n}f_{V_{i}}\left(\gamma_{i}^{n}V_{i}+\lambda V_{t}^{n}\right)\mathrm{d}t+\frac{1}{2}f_{SS}V_{t}^{n}\left(S\right)^{2}\mathrm{d}t+\sum_{i=1}^{n}f_{SV_{i}}\eta\rho V_{t}^{n}S\mathrm{d}t$
@@ 5. Implementation details
$u(x_{p}+y;\theta)=u(x_{p};\theta)+y-x_{p},\quad y>0.$
@@
Moreover, considering the moneyness $\frac{S}{K}$ instead of the stock price $S$, will return the relative option price $\frac{u}{K}$. Thus, we can approximate the option prices for all combinations of $S$ and $K$ with only one parameter.
@@
$|\quad Z^{l}$ $=\sigma_{1}\left(U^{z,l}\mathbf{x}+W^{z,l}X^{l}+b^{z,l}\right),\quad$ $l=1,\dots,L,$
$|\quad G^{l}$ $=\sigma_{1}\left(U^{g,l}\mathbf{x}+W^{g,l}X^{l}+b^{g,l}\right),\quad$ $l=1,\dots,L,$
$|\quad R^{l}$ $=\sigma_{1}\left(U^{r,l}\mathbf{x}+W^{r,l}X^{l}+b^{r,l}\right),\quad$ $l=1,\dots,L,$
$|\quad\!H^{l}$ $=\sigma_{1}\left(U^{h,l}\mathbf{x}+W^{h,l}\left(X^{l}\odot R^{l}\right)+b^{h,l}\right),\quad$ $l=1,\dots,L,$
@@
We consider a maturity of $T=1.0$ year, and take the number of time steps equal to $N=100$. We use $2000$ sampling stages in each time step, and in each sampling stage we take $600(n+1)$ samples, following the recommendations in Georgoulis et al. [28]. As for the sampling domain, we consider the moneyness $x\in[0.01,3.0]$, the Heston volatility $V\in[0.001,0.1]$ and the lifted Heston volatilities $V^{n,i}\in\left[-V_{i}^{\text{high}},V_{i}^{\text{high}}\right]$, with
@@
After sampling $V^{n,i}$ uniformly, we calculate $V^{n}$ and discard all combinations that result in a negative $V^{n}$. We set $x_{p}=2.0$ and use a learning rate $\alpha=3\times 10^{-4}$. The training was performed on the DelftBlue supercomputer [19], using a single NVidia Tesla V100S GPU .
@@ 6. Numerical results
$\left\|u_{t}-\nabla\cdot\left(A\nabla u\right)+\mathbf{b}\cdot\nabla u+ru\right\|_{L^{2}([0,T]\times\Omega)}^{2}+\left\|u(0,\mathbf{x})-\Psi(x)\right\|_{L^{2}(\Omega)}^{2}\to\min.$
@@
In order to have a fair comparison between the two methods, we use 200,000 sampling stages for the DGM. In each stage, we use $600(n+2)$ samples for the PDE in the interior domain and $600(n+1)$ samples for the initial and boundary conditions. The learning rate follows the schedule from Georgoulis et al. [28, eq. (24)].
@@ 6.1. Accuracy
In the Black–Scholes model, we only have one input variable, the asset price $S$. Figure 1 presents the relative $L^{2}$- and the absolute maximum error for the TDGF and the DGM methods in the Black–Scholes model. These errors are calculated over 41 evenly spaced grid points in the interval $[0.01,3]$. Both methods have a comparable (small) error.
@@
Figure 1. Errors of the two methods in the Black–Scholes model against time, with interest rate $r=0.05$ and volatility $\sigma=0.25$.
In the Heston model, we have two input variables, the asset price $S$ and the variance $V$. Figure 2 presents the option price against the moneyness level for different time instances, and Figure 3 presents the relative $L^{2}$- and the absolute maximum error for the TDGF and the DGM methods in the Heston model. These figures are presented for fixed $V_{0}=0.03$, however other values of $V_{0}$ between 0.01 and 0.09 provide similar results. The DGM is more accurate than the TDGF method, although the TDGF is just as accurate as in the Black–Scholes model, and the error is any case small, in the order of $10^{-3}$.
Figure 2. Relative European call option prices in the Heston model against the moneyness of the stock, compared to the COS and the DGM methods, for four different times to maturity with $r=0.0$ and $\eta=0.1$, $\rho=0.0$, $\kappa=0.01$, $V_{0}=0.03$ and $\lambda=2.0$.
@@
Figure 3. Error of the two methods in the Heston model against time, with $r=0.0$ and $\eta=0.1$, $\rho=0.0$, $\kappa=0.01$, $V_{0}=0.03$ and $\lambda=2.0$.
In the lifted Heston model, we have $n+1$ input variables, the asset price $S$ and the variances $V^{n,i}$, with $i=1,\dots,n$. Figure 4 presents the relative $L^{2}$- and the maximum absolute error for the TDGF and the DGM methods in this model, in the case $n=1$, using the same parameters as in the Heston model before. The numerical results are comparable for the two methods, and both methods even seem to perform slightly better than in the previous example.
@@
Figure 4. Errors of the two methods in the lifted Heston model with $n=1$ variance process against time, with $r=0.0$, $\eta=0.1$, $\rho=0.0$, $\kappa=0.01$, $V_{0}=0.01$ and $\lambda=2.0$.
Figure 5 presents the option price against the moneyness level, and Figure 6 presents the relative $L^{2}$- and the maximum absolute error for the TDGF and the DGM methods again in the lifted Heston model, for the same parameter set as previously, but now with $n=20$ variance factors. The errors of both methods are comparable and, although they have increased compared to the $n=1$ case, they are still low, in the order of $10^{-3}$.
Figure 5. Relative European call option prices in the lifted Heston model with $n=20$ variance processes against the moneyness of the stock, compared to the COS and the DGM methods, for four different times to maturity with $r=0.0$, $\eta=0.1$, $\rho=0.0$, $\kappa=0.01$, $V_{0}=0.01$ and $\lambda=2.0$.
@@
Figure 6. Errors of the two methods in the lifted Heston model with $n=20$ variance process against time, with $r=0.0$, $\eta=0.1$, $\rho=0.0$, $\kappa=0.01$, $V_{0}=0.01$ and $\lambda=2.0$.
Figure 7 presents the option price against the moneyness level, and Figure 8 presents the relative $L^{2}$- and the absolute maximum error for the two methods in the lifted Heston model, now for the parameter set suggested by Abi Jaber [1] with $n=1$. The main difference here is that the correlation between the asset price and the variance processes is no longer zero. The error in both methods has slightly increased with respect to the previous example, with the DGM performing slightly better than the TDGF. However, the overall error is still small, in the order of $10^{-3}$.
Figure 7. Relative European call option prices in the lifted Heston model with $n=1$ variance process against the moneyness of the stock, compared to the COS and the DGM methods, for four different times to maturity with $r=0.0$, $\eta=0.3$, $\rho=-0.7$, $\kappa=0.02$, $V_{0}=0.02$ and $\lambda=0.3$.
@@
Figure 8. Errors of the two methods in the lifted Heston model with $n=1$ variance process against time, with $r=0.0$, $\eta=0.3$, $\rho=-0.7$, $\kappa=0.02$, $V_{0}=0.02$ and $\lambda=0.3$.
Finally, Figure 9 presents the option price against the moneyness level, and Figure 10 presents the relative $L^{2}$- and the absolute maximum error for the two methods in the lifted Heston model, for the same parameter set as in the previous two figures, but now with $n=5$ and $\rho=-0.2$. The combination of increased dimension (i.e. 5 variance processes compared to one before) and decreased correlation results in the errors of both methods being similar to the previous case, and still small, in the order of $10^{-3}$. However, as the number of variance processes increase while the correlation is non-zero, then the errors in both methods are increasing rapidly, and we can no longer accommodate e.g. 20 variance factors as we did in the uncorrelated case.
Figure 9. Relative European call option prices in the lifted Heston model with $n=5$ variance processes against the moneyness of the stock, compared to the COS and the DGM methods, for four different times to maturity with $r=0.0$, $\eta=0.3$, $\rho=-0.2$, $\kappa=0.02$, $V_{0}=0.02$ and $\lambda=0.3$.
@@
Figure 10. Errors of the two methods in the lifted Heston model with $n=5$ variance processes against time, with $r=0.0$, $\eta=0.3$, $\rho=-0.2$, $\kappa=0.02$, $V_{0}=0.02$ and $\lambda=0.3$.

### v2

@@ 1. Introduction
Deep learning and neural network methods have also been applied in the field of rough stochastic volatility modeling. Horvath et al. [36] presented a neural network-based method that performs the calibration task for the full implied volatility surface in (rough) volatility models. Jacquier and Oumgari [37] introduced a deep learning–based algorithm to evaluate options in affine rough stochastic volatility models. Jacquier and Zuric [39] constructed a deep learning-based numerical algorithm to solve path-dependent partial differential equations arising in the context of rough volatility.
Many other methods have been utilized and applied to rough stochastic volatility models. Let us mention here, among others, that El Euch and Rosenbaum [22] have computed the characteristic function of the log‐price in the rough Heston model, thus opening the door for the application of Fourier transform methods for option pricing. Bayer et al. [11] proposed the application of adaptive sparse grids and quasi Monte Carlo methods for option pricing in the rough Bergomi model, while Bayer et al. [12] have applied Hairer’s theory of regularity structures in order to analyze rough volatility models. In addition, Bayer et al. [10], Horvath et al. [35] and Jacquier and Pannier [38], among others, have studied asymptotic properties of rough volatility models in various regimes.
The connection between partial differential equations (PDEs) and option pricing is deep and well-studied in the literature. Several methods have been developed in order to price options in diffusion or stochastic volatility models, especially in high-dimensional settings, using sparse grids, asymptotic expansions, finite difference and finite elements methods; see e.g. Reisinger and Wittum [46], Düring and Fournié [20], Hilber et al. [33] and the comprehensive book by Hilber et al. [34].
@@
The starting point for our research is the existence of a Markovian approximation for a rough volatility model or, more generally, another Markovian model that describes empirical data well; see e.g. Rømer [47] and Abi Jaber and Li [4]. Inspired by Georgoulis et al. [28], we consider the option pricing PDE, introduce a time-stepping discretization of the time derivative, and reformulate the spatial derivatives as an energy minimization problem; the main difference with Georgoulis et al. [28] is the presence of the drift (first order) term, which we treat in an explicit fashion in the discretization. This gradient flow formulation of the PDE gives rise to a suitable cost function for the training, via stochastic gradient descent, of deep neural networks that approximate the solution of the PDE. The main advantages of this method are twofold: on the one hand, the neural network from the previous time step is a good initial guess for the neural network optimization in the current time step, thus reducing the number of training stages necessary. On the other hand, the energy formulation allows to consider only first-order derivatives, thus reducing the training time in each time step / training stage significantly. In addition, we utilize available information about the qualitative behavior of option prices in order to assist the training of the neural network. More specifically, the architecture of the neural network respects the asymptotic behavior of option prices for large levels of moneyness, and adheres to a priori known bounds for option prices. The empirical results show that the proposed method is accurate, with a small overall error in the order of $10^{-3}$ and comparable with other deep learning methods, while the training times are significantly smaller compared to similar methods and do not grow significantly with the dimension of the underlying space.
@@ 2. Problem formulation
$\begin{cases}u_{t}(t,x)+\mathcal{A}u(t,x)+ru(t,x)=0,\quad&(t,x)\in(0,T]\times\Omega,\\ u(0,x)=\Psi(x),\quad&x\in\Omega,\end{cases}$ (2.1)
@@
Let us point out that, the order of the derivatives with respect to $x_{i}$ and $x_{j}$ in the second derivative term can be interchanged in (2.2), hence we can always choose $a^{ij}=a^{ji}$. Therefore, $A$ is a symmetric matrix.
@@ 3. Energy minimization and a deep gradient flow method
Let us now explain how we will develop our deep neural network method for solving the PDE (2.1). We proceed in three steps. First, we discretize the time derivative in Subsection 3.1 and introduce a time-stepping scheme. Second, we rewrite the solution of the PDE as an energy minimization problem in Subsection 3.2, which gives rise to a suitable cost function. Third, we approximate the minimizer by a neural network, which is trained using stochastic gradient descent, in Subsection 3.3
@@ 3.1. Time discretization
Let us divide the time interval $(0,T]$ into $N$ equally spaced intervals $(t_{k-1},t_{k}]$, with $h=t_{k}-t_{k-1}=\frac{1}{N}$ for $k=0,1,\dots,N$. Let $U^{k}$ denote the approximation to the solution of the PDE $u(t_{k})$ at time step $t_{k}$, using the following time discretization scheme
@@
where $F(u)=\mathbf{b}\cdot\nabla u$. Convergence analysis of this scheme appears in Akrivis et al. [5].
Remark 3.1.
Let us point out that non-uniform grids can also be used for the discretization of time, in order to employ a finer grid initially when the solution is less smooth. The numerical experiments in Section 4 show that this is not necessary in the examples we consider.
@@ 3.2. Variational formulation
We would like now to associate the discretized PDE in (3.1) with an energy functional, in order to define later a suitable cost function for the training of the neural network. In other words, we would like to find an energy functional $I(u)$ such that $U^{k}$ is a critical point of $I$. Let us rewrite (3.1) as follows:
@@
Assume that $A$ is symmetric and positive semi-definite. Then, $U^{k}$ solves (3.2) if and only if $U^{k}$ minimizes
@@
Let $v$ be a smooth function that is zero on the boundary. Then, for all such $v$, consider the function
@@
$\quad+\left.h\int_{\Omega}\frac{1}{2}\left(\left(\nabla v\right)^{\mathsf{T}}A\nabla\left(U^{k}+\tau v\right)+\left(\nabla\left(U^{k}+\tau v\right)\right)^{\mathsf{T}}A\nabla v\right)+r\left(U^{k}+\tau v\right)v+F\left(U^{k-1}\right)v\mathrm{d}x\right]_{\tau=0}$
@@
$=\int_{\Omega}\left(U^{k}-U^{k-1}\right)v+hA\nabla U^{k}\cdot\nabla v+h\left(rU^{k}+F\left(U^{k-1}\right)\right)v\mathrm{d}x$
@@
where in the second to last equality we used that $A$ is symmetric, and in the last equality we used the divergence theorem. This equality holds for all $v$ if and only if (3.2) holds. The second derivative is positive; indeed,
$\left(i^{k}\right)^{\prime\prime}(\tau)=(1+rh)\int_{\Omega}v^{2}\mathrm{d}x+h\int_{\Omega}\left(\nabla v\right)^{\mathsf{T}}A\nabla v\mathrm{d}x>0,$
since $A$ is positive semi-definite. Therefore, $\tau=0$ is indeed the minimizer. ∎
@@ 3.3. Neural network approximation
$L^{k}(\theta;\mathbf{x})$ $=\frac{|\Omega|}{2M}\sum_{i=1}^{M}\left(f^{k}(\mathbf{x}_{i};\theta)-f^{k-1}(\mathbf{x}_{i})\right)^{2}+hN^{k}\left(\theta;\mathbf{x}\right),$
@@
In order to minimize this cost function, we use a stochastic gradient descent-type algorithm, i.e. an iterative scheme of the form:
$\theta_{n+1}=\theta_{n}-\alpha_{n}\nabla_{\theta}L^{k}(\theta_{n};\mathbf{x});$
more specifically, we will use the Adam algorithm [41]. The hyper-parameter $\alpha_{n}$ is the step-size of our update, called the learning rate. An overview of the Time Deep Gradient Flow (TDGF) method appears in Algorithm 1.
@@ 4. Examples
Let us now outline the different models to which we will apply our method, and derive the appropriate form of the generator in the option pricing PDE for each of these models. In particular, we will treat the Black–Scholes model in Subsection 4.1, the Heston model in Subsection 4.2, and the lifted Heston model in Subsection 4.3. Let us point out that, straightforward but tedious computations, yield that the operator $A$ satisfies the assumptions of Lemma 3.2 for all three examples.
@@ 4.3. Lifted Heston model
$=f_{S}rS\mathrm{d}t-\sum_{i=1}^{n}f_{V_{i}}\left(\gamma_{i}^{n}V_{i}+\lambda V_{t}^{n}\right)\mathrm{d}t+\frac{1}{2}f_{SS}V_{t}^{n}S^{2}\mathrm{d}t+\sum_{i=1}^{n}f_{SV_{i}}\eta\rho V_{t}^{n}S\mathrm{d}t$
@@ 5. Implementation details
$u(x_{p}+y;\theta)=u(x_{p};\theta)+y,\quad y>0.$
@@
Moreover, we will consider the moneyness $\frac{S}{K}$ as an input variable instead of treating the asset price $S$ and the strike price $K$ separately, in order to reduce the size of the parameter space.
@@
$|\quad Z^{l}$ $=\sigma_{1}\left(U^{z,l}\mathbf{x}+W^{z,l}X^{l-1}+b^{z,l}\right),\quad$ $l=1,\dots,L,$
$|\quad G^{l}$ $=\sigma_{1}\left(U^{g,l}\mathbf{x}+W^{g,l}X^{l-1}+b^{g,l}\right),\quad$ $l=1,\dots,L,$
$|\quad R^{l}$ $=\sigma_{1}\left(U^{r,l}\mathbf{x}+W^{r,l}X^{l-1}+b^{r,l}\right),\quad$ $l=1,\dots,L,$
$|\quad\!H^{l}$ $=\sigma_{1}\left(U^{h,l}\mathbf{x}+W^{h,l}\left(X^{l-1}\odot R^{l}\right)+b^{h,l}\right),\quad$ $l=1,\dots,L,$
@@
We consider a maturity of $T=1.0$ year, and use the parameters sets for the Heston and the lifted Heston model from Heston [32, Table 1] and Abi Jaber and El Euch [3, page 321]. We set the number of time steps equal to $N=100$, use $2000$ sampling stages in each time step, while in each sampling stage we take 600 samples per dimension, i.e. $600(n+1)$ samples, following the recommendations in Georgoulis et al. [28]. Here, $n$ denotes the number of variance processes, in the case of stochastic volatility models. As for the sampling domain, we consider the moneyness $x\in[0.01,3.0]$, the Heston volatility $V\in[0.001,0.1]$ and the lifted Heston volatilities $V^{n,i}\in\left[-V_{i}^{\text{high}},V_{i}^{\text{high}}\right]$, with
@@
After sampling $V^{n,i}$ uniformly, we calculate $V^{n}$ and discard all combinations that result in a negative $V^{n}$. Moreover, we set $x_{p}=2.0$ In the optimization stage, we use the Adam algorithm [41] with a learning rate $\alpha=3\times 10^{-4}$, $(\beta_{1},\beta_{2})=(0.9,0.999)$ and zero weight decay. Finally, the training was performed on the DelftBlue supercomputer [19], using a single NVidia Tesla V100S GPU.
@@ 6. Numerical results
$\big{\|}u_{t}-\nabla\cdot\left(A\nabla u\right)+\mathbf{b}\cdot\nabla u+ru\big{\|}_{L^{2}([0,T]\times\Omega)}^{2}+\big{\|}u(0,\mathbf{x})-\Psi(x)\big{\|}_{L^{2}(\Omega)}^{2}\to\min.$
@@
In order to ensure a fair comparison between the two methods, we use 200,000 sampling stages for the DGM. In each stage, we use $600(n+2)$ samples for the PDE in the interior domain and $600(n+1)$ samples for the initial and boundary conditions. Moreover, we use the same network architecture as for the TDGF with the same number of layers and neurons per layer, and also use the Adam optimizer.
@@ 6.1. Accuracy
In the Black–Scholes model, we have only one input variable, the asset price $S$. Figure 6.1 presents the relative $L^{2}$- and the absolute maximum error for the TDGF and the DGM methods in the Black–Scholes model. These errors are calculated over 47 evenly spaced grid points in the interval $[0.01,3]$. Both methods have a comparable (small) error.
@@
Figure 6.1. Errors of the two methods in the Black–Scholes model against time, with interest rate $r=0.05$ and volatility $\sigma=0.25$.
In the Heston model, we have two input variables, the asset price $S$ and the variance $V$. Figure 6.2 presents the relative $L^{2}$- and the absolute maximum error for the TDGF and the DGM methods in the Heston model. These figures are presented for fixed $V_{0}=0.03$, however other values of $V_{0}$ between 0.01 and 0.09 provide similar results. The DGM is more accurate than the TDGF method, although the TDGF is just as accurate as in the Black–Scholes model, and the error is any case small, in the order of $10^{-3}$.
@@
Figure 6.2. Error of the two methods in the Heston model against time, with $r=0.0$ and $\eta=0.1$, $\rho=0.0$, $\kappa=0.01$, $V_{0}=0.03$ and $\lambda=2.0$.
In the lifted Heston model, we have $n+1$ input variables, the asset price $S$ and the variances $V^{n,i}$, with $i=1,\dots,n$. Figure 6.3 presents the relative $L^{2}$- and the maximum absolute error for the TDGF and the DGM methods in this model, in the case $n=1$, using the same parameters as in the Heston model before. The numerical results are comparable for the two methods, and both methods even seem to perform slightly better than in the previous example.
@@
Figure 6.3. Errors of the two methods in the lifted Heston model with $n=1$ variance process against time, with $r=0.0$, $\eta=0.1$, $\rho=0.0$, $\kappa=0.01$, $V_{0}=0.01$ and $\lambda=2.0$.
Figure 6.4 presents the relative $L^{2}$- and the maximum absolute error for the TDGF and the DGM methods again in the lifted Heston model, for the same parameter set as previously, but now with $n=20$ variance factors. The errors of both methods are comparable and, although they have increased compared to the $n=1$ case, they are still low, in the order of $10^{-3}$.
@@
Figure 6.4. Errors of the two methods in the lifted Heston model with $n=20$ variance process against time, with $r=0.0$, $\eta=0.1$, $\rho=0.0$, $\kappa=0.01$, $V_{0}=0.01$ and $\lambda=2.0$.
Figure 6.5 presents the relative $L^{2}$- and the absolute maximum error for the two methods in the lifted Heston model, now for the parameter set suggested by Abi Jaber [1] with $n=1$. The main difference here is that the correlation between the asset price and the variance processes is no longer zero. The error in both methods has slightly increased with respect to the previous example, with the DGM performing slightly better than the TDGF. However, the overall error is still small, in the order of $10^{-3}$.
@@
Figure 6.5. Errors of the two methods in the lifted Heston model with $n=1$ variance process against time, with $r=0.0$, $\eta=0.3$, $\rho=-0.7$, $\kappa=0.02$, $V_{0}=0.02$ and $\lambda=0.3$.
Finally, Figure 6.6 presents the relative $L^{2}$- and the absolute maximum error for the two methods in the lifted Heston model, for the same parameter set as in the previous two figures, but now with $n=5$ and $\rho=-0.2$. The combination of increased dimension (i.e. 5 variance processes compared to one before) and decreased correlation results in the errors of both methods being similar to the previous case, and still small, in the order of $10^{-3}$. However, as the number of variance processes increase while the correlation is non-zero, then the errors in both methods are increasing rapidly, and we can no longer accommodate e.g. 20 variance factors as we did in the uncorrelated case.
@@
Figure 6.6. Errors of the two methods in the lifted Heston model with $n=5$ variance processes against time, with $r=0.0$, $\eta=0.3$, $\rho=-0.2$, $\kappa=0.02$, $V_{0}=0.02$ and $\lambda=0.3$.